OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

I would just add the 2 functions together and the y's would cancel out. Then solve for x plug that back into one of the equations to get y.

OpenStudy (anonymous):

x - y = 10 ---------(i) 2x + y = 2 ---------(ii) rearranging (i) we get x = 10 + y substitute in (ii) 2(10 + y) + y = 2 20 + 2y + y = 2 3y = 2 - 20 y = - 18/3 y = -6 substitute in (i) x - (-6) = 10 x + 6 = 10 x = 10 - 6 x = 4 so solution is (4, -6) (ii) we used the substitution method because it was easy to get x in terms of y from (i) by a simple rearrangement of terms